OpenStudy (anonymous):

5 2/5 - 2 9/10? explain the steps. i will give you points :)

OpenStudy (mrdoe):

do you mean \[\frac{ 52 }{ 5 } - \frac{ 29 }{ 10 }\] If so just get a common denominator and subtract the numerators, then simplify

OpenStudy (anonymous):

can you explain the steps to find the answer? please :)

OpenStudy (mrdoe):

ex) \[2(\frac{ 52 }{ 5 })\] = \[\frac{ 104 }{ 10 }\] than because they share a common denominator we can subtract the numerators ignoring the denominator \[\frac{ 104 }{ 10} - \frac{ 29 }{ 10 } = \frac{ 75 }{ 10 } = \frac{ 15 }{ 2}\]

OpenStudy (help!!!!):

so first find the common denomenator for this question it is 10. so for 5 2/5 you multiply by 2 for the numerator and the denomator 2x2 2x5 which =5 4/10

OpenStudy (help!!!!):